# UFEval: Unified Fine-grained Evaluation with Task and Aspect Generalization

### v3

@@ Appendix C Appendix C: Experiment Detail
Expertise Expe (Li et al. 2023c) Academic QA Expertise refers to the model’s ability to accurately use domain-specific terminology when answering academic questions. This includes correctly applying terms relevant to the subject, such as those in science, history, or literature. For example, when asked about a historical event, the model should use terms like ’primary sources’ or ’historical context’ instead of vague words like ’thin’ or ’event’. Proper use of academic terms shows the model’s understanding of the subject.
Correctness Corr (Li et al. 2023c) Academic VQA The conciseness metric assesses whether the model provides clear, direct, and succinct answers to academic questions, avoiding unnecessary details. The accuracy metric evaluates if the responses correctly address the query based on the provided academic context, ensuring precise and relevant information is delivered.
Explainability Expl (Li et al. 2023c) Academic VQA Explainability is a criterion used to assess how clearly a model explains its reasoning when answering an academic question. It focuses on whether the model provides clear, logical steps and justifications for its answer, making it easy for users to understand how the answer was reached.
Interpretability Intp (Chen et al. 2023)
@@
Completion Comp (Chen et al. 2023)
@@
Riskiness Risk (Chen et al. 2023)
@@
Table 20: The definition and source of Task-specific Aspects. (Part 6)
@@
Alignment Ali (Xiong et al. 2024; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023; Xu et al. 2023a)
@@
Object Relationship ObjR (Li et al. 2024a; Huang et al. 2023)
@@
Spatial Alignment SpaA (Ghosh, Hajishirzi, and Schmidt 2023; Li et al. 2024a; Huang et al. 2023; Gokhale et al. 2022)
@@
Action Alignment ActA (Li et al. 2024a; Huang et al. 2023)
@@
Count Alignment CouA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023)
@@
Object Alignment ObjA (Li et al. 2024a; Ghosh, Hajishirzi, and Schmidt 2023)
@@
Attribute Alignment AttA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023)
@@
Color Alignment ColA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023)
@@
Table 21: The definition and source of Task-specific Aspects. (Part 7)
@@
Shape Alignment ShaA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023)
@@
Texture Alignment TexA (Li et al. 2024a; Huang et al. 2023; Ghosh, Hajishirzi, and Schmidt 2023)
@@
Advanced Alignment AdaA (Li et al. 2024a)
@@
Creativity Cre (Smilevski, Lalkovski, and Madjarov 2018; Liu et al. 2024a)
@@
Engagingness Eng (Smilevski, Lalkovski, and Madjarov 2018; Liu et al. 2024a)
@@
Table 22: The definition and source of Task-specific Aspects. (Part 8)
Aspect Abbr Source Sub-Task Definition
Completeness Com (Smilevski, Lalkovski, and Madjarov 2018; Liu et al. 2024a)
@@
Completeness Com (Liu et al. 2024a; Koupaee and Wang 2018)
Multimodal Script
Generation
Feasibility Fea (Liu et al. 2024a; Koupaee and Wang 2018)
Multimodal Script
Generation
Helpfulness Hel (Liu et al. 2024a; Koupaee and Wang 2018)
Multimodal Script
Generation
Safety Saf (Liu et al. 2024a; Koupaee and Wang 2018)
Multimodal Script
Generation
Reasonableness Reas (Liu et al. 2024a)
Storytelling
Generation
Fulfillment Full (Liu et al. 2024a)
Storytelling
Generation
Originality Orig (Liu et al. 2024a)
Activitynet
Continuation
Simplification Simp (Liu et al. 2024a)
Activitynet
Continuation
Table 23: The definition and source of Task-specific Aspects. (Part 9)
@@
Table 24: Aspects used by evaluators in the meta-benchmark evaluations. (Part 1)
@@
Table 25: Aspects used by evaluators in the meta-benchmark evaluations. (Part 2)
@@
GenAI-Bench & Pickapic - Alignment Alignment between text and image is evaluated based on how well the image matches the details in the text. This includes the correct objects, their attributes (e.g., color, size), spatial relationships (e.g., left, right), and actions. The alignment should consider multiple dimensions. For example, if the text describes “two boys wearing white clothes, one on the left and one on the right, playing badminton,” the image should show two boys in white clothes, placed correctly in the scene, with a badminton game happening between them.
@@
Table 27: Aspects used by evaluators in the meta-benchmark evaluations. (Part 4)
Appendix D Appendix D: More Experiment Results
In-domain Evaluation
Experimental Results for FRAUAs-OOD
Specific Experimental Results
We report the specific average accuracy of each evaluator for every aspect within each task in Figure 11 and Figure 3 of the main text. The alignment is measured against: (1) GPT-4o on the FRA-ID and FRA-OOD, and (2) human annotators on the FRA-ID-H and FRA-OOD-H. Detailed results are presented in Tables 30 to 37. The reported average accuracy is computed by aggregating the same aspect across different sub-tasks within the same task. We also include the comprehensive results on MLLM-as-a-Judge in Table 46. The specific accuracy of multi-aspect assessment learning is show in Table 52 and  54.
Experiment Results of UFEval-72B
We also train UFEval-72B using the same fine-tuning method as UFEval-7B with Qwen2-VL-72B-Instruct as the backbone and test it under the same experiment settings. The results, shown in Figures 14 and 15 and Tables 47 to 49, indicate that UFEval-72B offers some improvement over UFEval-7B.
Appendix E Appendix E: Human Annotation
The user interface used for human annotation is shown in Figures 16 and 17. For each aspect, we adopt a 5-point Likert scale. Additionally, to account for cases where certain instructions are not applicable to TAs, we include a "-1" option to indicate inapplicability. To maintain high-quality annotation, we engage three experts to annotate each data point concurrently. These experts, recruited from graduate students to enhance annotation reliability, are provided with guidance whenever they encounter challenges during the annotation process, enabling them to focus on multiple aspects such as helpfulness, clarity, comprehensiveness, and more.
Appendix F Appendix F: The Prompt Template
Appendix G Appendix G: Qualitative Examples
Table 64, Table 65, Table 66, Tables 67 to 69 show some good qualitative examples of feedback generated by UFEval. Table 70, Table 71, and Table 72 shows a relatively inferior example of feedback generated by UFEval.
Method GPT-4o Alignment
NLG IU ITIG
Claude-3.5 77.3 77.6 83.0
Qwen2VL-72B 69.7 71.4 65.3
Qwen2VL-7B 49.9 56.5 44.4
Themis 54.2 - -
LLaVA-Critic - 50.0 -
Ours 82.8 83.1 90.4
Table 28: Average accuracy on FRAUAs-OOD.
@@ Appendix G Appendix G: Qualitative Examples
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 50.0 43.8 28.1 34.4 51.0
Alignment Ali 67.5 70.0 47.5 45.0 55.0
Attribute Alignment AttA 62.5 62.5 52.5 37.5 65.0
Color Alignment ColA 65.0 60.0 57.5 37.5 49.5
Count Alignment CouA 55.0 60.0 50.0 20.0 60.0
Fidelity Fid 30.0 42.5 40.0 33.0 35.0
Harmfulness Har 100 62.5 95.0 82.5 95.0
Object Alignment ObjA 52.5 57.5 50.0 27.5 57.5
Object Relationship ObjR 60.0 50.0 45.0 32.5 52.5
Shape Alignment ShaA 65.0 55.0 60.0 35.0 70.0
Spatial Alignment SpaA 57.5 60.0 37.5 27.5 57.5
Texture Alignment TexA 52.5 60.0 50.0 50.0 62.5
@@
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 62.7 56.6 51.8 42.2 50.6
Alignment Ali 75.8 61.2 55.1 41.0 65.2
Attribute Alignment AttA 75.2 62.1 46.6 49.7 62.1
Color Alignment ColA 80.7 68.3 62.1 47.2 75.8
Count Alignment CouA 68.5 74.1 48.1 27.8 66.7
Fidelity Fid 71.4 51.1 49.3 44.9 54.2
Harmfulness Har 94.3 80.6 89.0 77.5 82.8
Object Alignment ObjA 64.8 68.5 46.3 27.8 53.7
Object Relationship ObjR 72.0 62.2 35.7 26.6 63.6
Shape Alignment ShaA 72.0 50.3 46.6 48.4 55.3
Spatial Alignment SpaA 68.4 54.1 34.7 34.7 53.1
Texture Alignment TexA 75.2 65.2 39.1 43.5 59
Comparison Cmpa 77.5 80.0 50.0 37.5 70
Negation Nega 75.6 61.0 53.7 46.3 68.3
Scene Alignment ScnA 71.4 57.1 49.1 58.0 58.9
Universal Univ 72.5 62.5 62.5 42.5 67.5
@@
Table 46: Comprehensive results on MLLM-as-a-Judge. *: The results for GPT-4V and LLaVA-Critic are taken from the original papers (Chen et al. 2024) and (Xiong et al. 2024), respectively. All other results are obtained from our experiments using their publicly available codebases.
Method GenAI-Bench Winoground Pickapic
diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$)
GPT-4o 55.6 69.5 62.6 73.0 73.0 69.5 54.4 59.2
Claude-3.5 55.6 71.0 71.2 73.0 69.2 71.1 49.1 53.7
Qwen2VL-72B 49.1 52.6 46.7 60.9 57.6 55.0 38.6 38.3
Qwen2VL-7B 35.8 38.0 33.4 42.5 46.1 40.6 38.1 40.2
VisionReward 51.0 66.4 60.2 64.1 74.9 66.4 48.9 58.0
ImageReward 48.6 64.9 54.0 58.2 69.2 60.4 48.8 55.7
Ours 53.6 65.5 57.5 59.1 80.7 65.7 50.0 57.3
Ours-72B 55.5 69.0 61.5 63.4 62.5 62.4 52.0 58.6
Table 47: Evaluation as LLM-as-a-Judge for IG task. We evaluate baselines across three meta-benchmarks.
Method SummEval MANS MT-Bench
GPT-4o 58.0 64.2 79.1 85.1 64.3 72.8 60.1 67.1 65.3 72.3 68.5 70.9 83.5
Themis 60.7 62.1 81.8 86.0 73.3 77.7 54.4 54.6 67.5 70.1 44.2 43.6 37.7
Auto-J - - - - - - - - 60.4 67.0 68.2 73.0 85.5
Ours 64.6 71.8 75.2 83.5 74.7 84.5 61.3 67.6 69.0 76.9 69.3 74.9 88.3
Table 48: Evaluation as LLM-as-a-Judge for NLG task. We evaluate baselines across three meta-benchmarks.
Method WildVision MLLM-as-a-Judge VLRewardBench
diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$)
GPT-4o 55.3 70.1 73.3 58.1 67.0 50.2 81.4 74.8 68.8
Claude-3.5 53.3 67.3 61.2 58.4 68.3 38.5 82.6 66.1 62.4
Qwen2VL-72B 50.3 59.6 65.5 54.6 58.5 50.8 75.4 70.7 65.6
Qwen2VL-7B 39.2 40.6 23.1 41.3 44.6 45.1 62.8 62.5 56.8
LLaVA-Critic 53.0 66.0 59.6 55.6 65.5 42.0 41.2 60.0 47.7
Ours 53.9 68.6 66.5 57.2 67.0 46.4 57.7 71.1 58.4
Ours-72B 54.5 68.6 69.9 58.2 73.1 44.3 61.6 72.1 59.3
Table 49: Evaluation as LLM-as-a-Judge for IU tasks. We evaluate baselines across three meta-benchmarks.
Annotators NLG IU IG ITIG
human 1-2 79.0 85.6 76.8 95.7
human 1-3 87.2 90.2 80.8 91.4
human 2-3 91.6 95.3 92.7 95.6
Table 50: Inter Annotator Agreement of each annotator on the FRA-ID-H.
Annotators NLG IU IG ITIG
human 1-2 80.3 95.0 76.1 84.7
human 1-3 88.1 81.3 73.9 82.4
human 2-3 77.7 79.1 77.4 78.6

### v4

@@ C.3 Aspects Used in the public benchmark
Expertise Expe (Li et al., 2023c) Academic QA Expertise refers to the model’s ability to accurately use domain-specific terminology when answering academic questions. This includes correctly applying terms relevant to the subject, such as those in science, history, or literature. For example, when asked about a historical event, the model should use terms like ’primary sources’ or ’historical context’ instead of vague words like ’thin’ or ’event’. Proper use of academic terms shows the model’s understanding of the subject.
Correctness Corr (Li et al., 2023c) Academic VQA The conciseness metric assesses whether the model provides clear, direct, and succinct answers to academic questions, avoiding unnecessary details. The accuracy metric evaluates if the responses correctly address the query based on the provided academic context, ensuring precise and relevant information is delivered.
Explainability Expl (Li et al., 2023c) Academic VQA Explainability is a criterion used to assess how clearly a model explains its reasoning when answering an academic question. It focuses on whether the model provides clear, logical steps and justifications for its answer, making it easy for users to understand how the answer was reached.
Interpretability Intp (Chen et al., 2023)
@@
Completion Comp (Chen et al., 2023)
@@
Riskiness Risk (Chen et al., 2023)
@@
Originality Orig (Liu et al., 2024a)
Activitynet
Continuation
Simplification Simp (Liu et al., 2024a)
Activitynet
Continuation
Table 21: The definition and source of Task-specific Aspects (TAs). (Part 6)
@@
Fulfillment Full (Liu et al., 2024a)
Storytelling
Generation
Alignment Ali (Xiong et al., 2024; Huang et al., 2023; Ghosh et al., 2023; Xu et al., 2023a)
@@
Object Relationship ObjR (Li et al., 2024a; Huang et al., 2023)
@@
Spatial Alignment SpaA (Ghosh et al., 2023; Li et al., 2024a; Huang et al., 2023; Gokhale et al., 2022)
@@
Action Alignment ActA (Li et al., 2024a; Huang et al., 2023)
@@
Count Alignment CouA (Li et al., 2024a; Huang et al., 2023; Ghosh et al., 2023)
@@
Object Alignment ObjA (Li et al., 2024a; Ghosh et al., 2023)
@@
Attribute Alignment AttA (Li et al., 2024a; Huang et al., 2023; Ghosh et al., 2023)
@@
Color Alignment ColA (Li et al., 2024a; Huang et al., 2023; Ghosh et al., 2023)
@@
Completeness Com (Liu et al., 2024a; Koupaee & Wang, 2018)
Multimodal Script
Generation
Feasibility Fea (Liu et al., 2024a; Koupaee & Wang, 2018)
Multimodal Script
Generation
Helpfulness Hel (Liu et al., 2024a; Koupaee & Wang, 2018)
Multimodal Script
Generation
Table 22: The definition and source of Task-specific Aspects (TAs). (Part 7)
@@
Safety Saf (Liu et al., 2024a; Koupaee & Wang, 2018)
Multimodal Script
Generation
Shape Alignment ShaA (Li et al., 2024a; Huang et al., 2023; Ghosh et al., 2023)
@@
Texture Alignment TexA (Li et al., 2024a; Huang et al., 2023; Ghosh et al., 2023)
@@
Advanced Alignment AdaA (Li et al., 2024a)
@@
Creativity Cre (Smilevski et al., 2018; Liu et al., 2024a)
@@
Engagingness Eng (Smilevski et al., 2018; Liu et al., 2024a)
@@
Completeness Com (Smilevski et al., 2018; Liu et al., 2024a)
@@
Table 23: Aspects used by evaluators in the benchmark evaluations. (Part 1)
@@
Table 24: Aspects used by evaluators in the benchmark evaluations. (Part 2)
@@
Table 26: Aspects used by evaluators in the benchmark evaluations. (Part 4)
@@
GenAI-Bench & Pick-a-Pic - Alignment Alignment between text and image is evaluated based on how well the image matches the details in the text. This includes the correct objects, their attributes (e.g., color, size), spatial relationships (e.g., left, right), and actions. The alignment should consider multiple dimensions. For example, if the text describes “two boys wearing white clothes, one on the left and one on the right, playing badminton,” the image should show two boys in white clothes, placed correctly in the scene, with a badminton game happening between them.
@@
Appendix D More Experiment Results
D.1 In-domain Evaluation
@@
D.2 Experimental Results for FRAUAs-OOD
Table 28: The results of average accuracy on FRAUAs-OOD for UFEval.
Method GPT-4o Alignment
NLG IU ITIG
Claude-3.5 77.3 77.6 83.0
Qwen2VL-72B 69.7 71.4 65.3
Qwen2VL-7B 49.9 56.5 44.4
Themis 54.2 - -
LLaVA-Critic - 50.0 -
Ours 82.8 83.1 90.4
D.3 Experiment Results of UFEval-72B
We also train UFEval-72B using the same fine-tuning method as UFEval-7B with Qwen2-VL-72B-Instruct as the backbone and test it under the same experiment settings. The results, shown in Figures 14 and 15 and Tables 29 to 31, indicate that UFEval-72B offers some improvement over UFEval-7B.
Table 29: Evaluation as MLLM-as-a-Judge for IG task. We evaluate baselines across three benchmarks.
Method GenAI-Bench Winoground Pick-a-Pic
diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$)
GPT-4o 55.6 69.5 62.6 73.0 73.0 69.5 54.4 59.2
Claude-3.5 55.6 71.0 71.2 73.0 69.2 71.1 49.1 53.7
Qwen2VL-72B 49.1 52.6 46.7 60.9 57.6 55.0 38.6 38.3
Qwen2VL-7B 35.8 38.0 33.4 42.5 46.1 40.6 38.1 40.2
VisionReward 51.0 66.4 60.2 64.1 74.9 66.4 48.9 58.0
ImageReward 48.6 64.9 54.0 58.2 69.2 60.4 48.8 55.7
Ours 53.6 65.5 57.5 59.1 80.7 65.7 50.0 57.3
Ours-72B 55.5 69.0 61.5 63.4 62.5 62.4 52.0 58.6
Table 30: Evaluation as MLLM-as-a-Judge for IU tasks. We evaluate baselines across three benchmarks.
Method WildVision MLLM-as-a-Judge VLRewardBench
diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$)
GPT-4o 55.3 70.1 73.3 58.1 67.0 50.2 81.4 74.8 68.8
Claude-3.5 53.3 67.3 61.2 58.4 68.3 38.5 82.6 66.1 62.4
Qwen2VL-72B 50.3 59.6 65.5 54.6 58.5 50.8 75.4 70.7 65.6
Qwen2VL-7B 39.2 40.6 23.1 41.3 44.6 45.1 62.8 62.5 56.8
LLaVA-Critic 53.0 66.0 59.6 55.6 65.5 42.0 41.2 60.0 47.7
Ours 53.9 68.6 66.5 57.2 67.0 46.4 57.7 71.1 58.4
Ours-72B 54.5 68.6 69.9 58.2 73.1 44.3 61.6 72.1 59.3
Table 31: Evaluation as MLLM-as-a-Judge for NLG task.
Method SummEval MANS MT-Bench
GPT-4o 58.0 64.2 79.1 85.1 64.3 72.8 60.1 67.1 65.3 72.3 68.5 70.9 83.5
Themis 60.7 62.1 81.8 86.0 73.3 77.7 54.4 54.6 67.5 70.1 44.2 43.6 37.7
Auto-J - - - - - - - - 60.4 67.0 68.2 73.0 85.5
Ours 64.6 71.8 75.2 83.5 74.7 84.5 61.3 67.6 69.0 76.9 69.3 74.9 88.3
D.4 Specific Experimental Results
We report the specific average accuracy of each evaluator for every aspect within each task in Figure 11 and Figure 3 of the main text. The alignment is measured against: (1) GPT-4o on the FRA-ID and FRA-OOD, and (2) human annotators on the FRA-ID-H and FRA-OOD-H. Detailed results are presented in Tables 32 to 47. The reported average accuracy is computed by aggregating the same aspect across different sub-tasks within the same task. We also include the comprehensive results on MLLM-as-a-Judge in Table 48. The specific accuracy of multi-aspect assessment learning is show in Table 49 and  51.
@@ D.4 Specific Experimental Results
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 50.0 43.8 28.1 34.4 51.0
Alignment Ali 67.5 70.0 47.5 45.0 55.0
Attribute Alignment AttA 62.5 62.5 52.5 37.5 65.0
Color Alignment ColA 65.0 60.0 57.5 37.5 49.5
Count Alignment CouA 55.0 60.0 50.0 20.0 60.0
Fidelity Fid 30.0 42.5 40.0 33.0 35.0
Harmfulness Har 100 62.5 95.0 82.5 95.0
Object Alignment ObjA 52.5 57.5 50.0 27.5 57.5
Object Relationship ObjR 60.0 50.0 45.0 32.5 52.5
Shape Alignment ShaA 65.0 55.0 60.0 35.0 70.0
Spatial Alignment SpaA 57.5 60.0 37.5 27.5 57.5
Texture Alignment TexA 52.5 60.0 50.0 50.0 62.5
@@
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 62.7 56.6 51.8 42.2 50.6
Alignment Ali 75.8 61.2 55.1 41.0 65.2
Attribute Alignment AttA 75.2 62.1 46.6 49.7 62.1
Color Alignment ColA 80.7 68.3 62.1 47.2 75.8
Count Alignment CouA 68.5 74.1 48.1 27.8 66.7
Fidelity Fid 71.4 51.1 49.3 44.9 54.2
Harmfulness Har 94.3 80.6 89.0 77.5 82.8
Object Alignment ObjA 64.8 68.5 46.3 27.8 53.7
Object Relationship ObjR 72.0 62.2 35.7 26.6 63.6
Shape Alignment ShaA 72.0 50.3 46.6 48.4 55.3
Spatial Alignment SpaA 68.4 54.1 34.7 34.7 53.1
Texture Alignment TexA 75.2 65.2 39.1 43.5 59
Comparison Cmpa 77.5 80.0 50.0 37.5 70
Negation Nega 75.6 61.0 53.7 46.3 68.3
Scene Alignment ScnA 71.4 57.1 49.1 58.0 58.9
Universal Univ 72.5 62.5 62.5 42.5 67.5
@@
Table 48: Comprehensive results on MLLM-as-a-Judge. *: The results for GPT-4V and LLaVA-Critic are taken from the original papers (Chen et al., 2024) and (Xiong et al., 2024), respectively. All other results are obtained from our experiments using their publicly available codebases.
@@
Appendix E Human Annotation
The user interface used for human annotation is shown in Figures 16 and 17. For each aspect, we adopt a 5-point Likert scale. Additionally, to account for cases where certain instructions are not applicable to TAs, we include a "-1" option to indicate inapplicability. To maintain high-quality annotation, we engage three humans to annotate each data point concurrently. These humans, recruited from graduate students to enhance annotation reliability, are provided with guidance whenever they encounter challenges during the annotation process, enabling them to focus on multiple aspects such as helpfulness, clarity, comprehensiveness, and more.